OpenStudy (taylev105):

Convert 7.54545454... to a rational number in the form of a /b , where b ≠ 0.

OpenStudy (taylev105):

will medal and fan

OpenStudy (anonymous):

let x=7.545454...(1) multiply by 100 then subtract (1) from it. calculate x.

OpenStudy (taylev105):

753?

OpenStudy (taylev105):

idk what X is im clueless I hate q like this

OpenStudy (anonymous):

100 x=754.5454.... 99 x=745 x=?

OpenStudy (anonymous):

correction it is 747 not 745

OpenStudy (taylev105):

where did the 99 come from ?

OpenStudy (anonymous):

100 x-x=99 x

OpenStudy (anonymous):



OpenStudy (taylev105):

ok so 7 54/99 is the answer I got a private tutor

OpenStudy (anonymous):

\[\frac{ 747 }{ 99 }\]

OpenStudy (anonymous):



OpenStudy (taylev105):

yep